# Priority aware grouping-based multihop routing scheme for RIS-assisted wireless networks

### v1

@@ Abstract
Reconfigurable intelligent surfaces (RISs) is a novel communication technology that has been recognized and recently presented as a candidate for beyond fifth generation wireless communication technology. In this paper, we propose a priority aware user traffic dependent grouping based multihop routing scheme for a RIS-assisted millimeter wave (mmWave) device-to-device (D2D) communication network with spatially correlated channels. Specifically, the proposed scheme exploits the priority of the users (based on their respective delay constrained applications) and the aspect of spatial correlation in the narrowly spaced reflecting elements of the RISs. In this context, we establish a multihop connection for information transfer from one of the users to its desired receiver based on the other users in the neighbourhood, their respective traffic characteristics, and the already deployed RISs in the surroundings. Moreover, we also take into account the impact of considering practical discrete phase shifts at the RIS patches instead of its ideal continuous counterpart. Furthermore, we claim as well as demonstrate that the existing classic least remaining distance (LRD)-based approach is not always the optimal solution. Finally, numerical results demonstrate the advantages of the proposed strategy and that it significantly outperforms the existing benchmark schemes in terms of system performance metrics such as data throughput, energy consumption, as well as energy efficiency.
@@ I Introduction
In recent years, due to applications such as the Internet of Things (IoT) and enhanced mobile broadband communication, wireless traffic has increased exponentially; it is projected to expand by more than five times in the $2023-2028$ period [1]. As a result, various technologies such as adaptive modulation, beamforming, and multiple access frameworks have been developed in the last few decades to support this overwhelming data traffic. The unifying theme behind all these advancements is to intelligently exploit and adapt to the randomly fluctuating wireless channel instead of having a control over the same. On the contrary, a new technology termed as the reconfigurable intelligent surfaces (RISs) is aimed at controlling the random nature of the channel for the purpose of its maximum utilization [2]. RISs essentially consist of arrays of reconfigurable passive elements embedded on a flat metasurface, which effectively ‘controls’ the channel instead of adapting to its varying nature [3]. These passive elements can be switched ON and OFF by adjusting the bias voltage. Moreover, without the need of any radio-frequency chains, they can reflect the incident signal in a desired direction. This results in reducing the implementation cost, which also enhances the system energy efficiency [4].
@@
In this context, RIS-assisted device-to-device (D2D) communication [5, 6, 7] forms a very relevant and interesting direction of research. The authors in [5] focus on uplink D2D-enabled cellular networks assisted by an RIS, where multiple D2D links share the same spectrum. In [6], the outage and rate performance of an RIS-assisted D2D communication underlaying cellular network is investigated by considering a generalized Nakagami-$m$ fading scenario. By using the concept of Markov service processes, the work in [7] analyzes the throughput of a delay constrained RIS-assisted D2D communication scenario. Apart from system energy efficiency improvement by using RISs, the aspect of obtaining a high data rate can also be addressed by the usage of high frequency signals, for example, the millimeter waves (mmWaves) [8]. Note that, although mmWaves solve the problem of enhanced data rate, it has its own set of shortcomings such as significant propagation and high penetration losses. As a result, RIS assisted mmWave-based D2D network appears to be the solution for such scenarios, especially where the direct line of sight (LoS) link is of inadequate strength. Note that, to fully exploit the advantages of such networks, the aspect of strategic RIS placement is extremely crucial as they need to have clear LoS links with both the devices that intend to communicate with each other. Optimal RIS placement [9, 10, 11] can significantly extend the coverage of such networks with a minimum number of RISs, which effectly reduces both the hardware and installation expenditure.
@@
We further note that the RIS is essentially a passive device that simply reflects the incoming signal as well as the noise to its desired direction by tuning its parameter. Moreover, any given pair of users always communicate with each other for only a finite amount of time. Thus, providing an RIS for each communicating pair, may result in unnecessary wastage of resources (deployment of a large number of RISs). A possible solution to this problem is to adopt a cooperative multihop approach [12, 13, 14], i.e., apart from the RISs, the other intermediate users (IUs) present in the surroundings, depending on whether they are busy or idle, can be used as relays, namely, decode and forward (DF) or amplify and forward (AF). While AF relays have low processing cost, DF relays have high reliability as they forward not the entire received information-plus-noise mixture but only the information portion to the next hop [15]. Thus, the option of having the idle surrounding users to act as DF relays seems beneficial. However, the role of RISs cannot be totally overlooked as in case of delay constrained scenarios and/or idle IU unavailability, they provide the aspect of guaranteed reliability [16].
@@
Moreover, as far as the structure of a RIS is concerned, there are certain aspects that are critical to be considered and analyzed in this multihop framework. Firstly, the arrays of the reconfigurable elements are very narrowly spaced, which implies that the channels between the consecutive reflecting elements and the source/destination are spatially correlated [17]. Hence, to avoid the huge channel estimation overhead in RIS-based systems, the works in [18, 19, 20] propose grouping strategies where the nearby RIS elements are grouped together into smaller non-overlapping surfaces with identical phase shifts. Secondly, most existing works assume a continuous nature of the phase shifts that are offered by a RIS whereas in practical systems, they are actually discrete in nature [21]. Hence, these limited phase shifts do have a significant impact on the system performance. These aspects of a RIS cannot be overlooked if we are interested in investigating the practically feasible of a joint user and RIS based multihop framework. Furthermore, least remaining distance (LRD)-based approach is a standard procedure in literature to establish a multihop connection between two users, whether the sole objective is to reduce the effective distance between the two in each hop [22]. But, unlike in this case, this approach stands valid if and only if we do not take the channel conditions and the availability of the IUs into consideration.
@@ II-A Network Topology
Each RIS has $\mathcal{K}$ reflecting elements, which are effectively controlled to adjust both the amplitude and phase of the incident waveform  [^1] Here we do not consider secondary reflections from the RISs. However, the proposed framework is general and it can be readily extended to the case where the aspect of secondary reflection is incorporated [24].. However, for the sake of simplicity and mathematical tractability, the amplitude factor is set to unity and it is only the phase that is tuned or optimized [25]. To reduce the channel estimation overhead, we employ a grouping strategy, where nearby $K_{g}$ reflecting elements are grouped together, such that all the elements belonging to the same group introduce identical phase shift [18]. Specifically, the $r$-th RIS $R_{r}$, consisting of $\mathcal{K}$ reflecting elements, is subdivided into $B$ non-overlapping surfaces $R_{r,b}$ $\forall$ $1\leq b\leq B$ such that $BK_{g}=\mathcal{K}$, where the number of partitions and the resulting sub-surfaces are determined a-priori [20]. Without any loss of generality, we consider that each group has an equal number of elements and a particular group can serve one service request at a time [26]. In this scenario, $R_{r,b}$ has two possible states: ON and OFF. If an element is in ON state, the phase of an incident signal can be changed to a desired direction and it cannot reflect while in OFF state. Finally, an example of the considered scenario is demonstrated in Fig. 1, where $S_{1}$ has the highest priority and the elements of the RIS are grouped into four non-overlapping surfaces, and the signal from $S_{1}$ reaches $D_{1}$ by using the path $S_{1}-U_{1}-R_{1}-U_{3}-D_{1}$.
@@ II-B Channel Model
Here $K_{p,q}$ is the Rician factor corresponding to the channel gain $h_{p,q}$ and $I_{0}(\cdot)$ denotes the zero-order modified Bessel function of the first kind. Moreover, note that direct $S-D$ path does not exist, i.e., as per our assumption, $h_{S_{i},D_{i}}=0$ $\forall$ $i$.
@@ III Proposed strategy
In this section, we propose a priority aware scheduling strategy for grouping-based RIS-assisted wireless networks. Here, we consider a delay-constrained scenario, where the data from $S_{i}$ must arrive at $D_{i}$ within the delay bound $T_{d_{i}}$. Firstly, we characterize the achievable data rate for a grouping-based RIS-aided scenario. Secondly, based on the channel condition and the number of coding bits being used at the RIS, we comment on the suitable group selection strategy. Thirdly, depending on the channel condition as well as the delay sensitivity of the application, a novel scheduling strategy is proposed for information transfer between any two IUs. Note that, we also comment on the appropriate choice of modulation scheme and appropriate next hop selection criteria. Finally, we provide an illustrative example of the proposed framework.
@@ III-A Achievable Rate Characterization of Grouping-based RIS
$h_{S_{i}/U_{j},R_{r,b}}$ $=\left(h_{I1}\sum\limits_{i=1}^{K_{g}}\mu_{1,i}+\sum\limits_{i=1}^{K_{g}}h_{Ii}\sqrt{1-\mu_{1,i}^{2}}\right)$
$+j\left(h_{Q1}\sum\limits_{i=1}^{K_{g}}\mu_{1,i}+\sum\limits_{i=1}^{K_{g}}h_{Qi}\sqrt{1-\mu_{1,i}^{2}}\right)$ (7)
@@
where $\zeta=\dfrac{P\rho_{\rm L}^{2}\left(d_{U_{j},R_{r}}d_{R_{r},U_{k}}\right)^{-\alpha}}{\sigma_{0}^{2}}$, $\phi_{r,b}$ is the associated phase shift provided at $R_{r,b}$, $T_{c}$ is the channel coherence time normalized to the number of slots, and $\sigma_{0}$ is the power of the circularly symmetric zero mean additive white Gaussian noise (AWGN) at $U_{k}$. Moreover, the quantity $\dfrac{2}{T_{c}}$ denotes the feedback and processing delay. Note that unlike the conventional RIS-based approach, $\gamma_{j,k}^{r,b}$ does not involve a diagonal phase shift matrix of $K_{g}$ non-zero elements. On the contrary, all the elements of $R_{r,b}$ provide a common phase shift $\phi_{r,b}$ to the incoming signal. Hence, by rewriting the complex composite channels as $h_{U_{j},R_{r,b}}\!=\!\Big{|}h_{U_{j},R_{r,b}}\Big{|}e^{-j\theta_{h}}$ and $g_{R_{r,b},U_{k}}\!=\!\Big{|}g_{R_{r,b},U_{k}}\Big{|}e^{-j\theta_{g}}$, we obtain the optimal phase shift corresponding to $R_{r,b}$ as $\phi_{r,b}^{\rm opt}=\theta_{h}+\theta_{g}$ such that the maximum attainable SNR is given by (III-A).
@@
=$\overset{(a)}{=}\frac{P\rho_{\rm L}^{2}\left(d_{U_{j},R_{r}}d_{R_{r},U_{k}}\right)^{-\alpha}}{\sigma_{0}^{2}}\left(\left(h_{I1}\sum\limits_{i=1}^{K_{g}}\mu_{1,i}+\sum\limits_{i=1}^{K_{g}}h_{Ii}\sqrt{1-\mu_{1,i}^{2}}\right)^{2}+\left(h_{Q1}\sum\limits_{i=1}^{K_{g}}\mu_{1,i}+\sum\limits_{i=1}^{K_{g}}h_{Qi}\sqrt{1-\mu_{1,i}^{2}}\right)^{2}\right)$
×$\times\left(\left(g_{I1}\sum\limits_{i=1}^{K_{g}}\mu_{1,i}+\sum\limits_{i=1}^{K_{g}}g_{Ii}\sqrt{1-\mu_{1,i}^{2}}\right)^{2}+\left(g_{Q1}\sum\limits_{i=1}^{K_{g}}\mu_{1,i}+\sum\limits_{i=1}^{K_{g}}g_{Qi}\sqrt{1-\mu_{1,i}^{2}}\right)^{2}\right),$ (10)
@@ III-B Group Selection Criteria at an RIS
$\!\!h_{U_{j},R_{r,b}}\!\!$ $=\Bigg{(}h_{I_{1}}\sum_{i=1}^{K_{g}}\mu_{1,i}+h_{I_{1}}\sum_{i=1}^{K_{g}}\sqrt{1-\mu^{2}_{1,i}}\Bigg{)}$
$+j\Bigg{(}h_{Q_{1}}\sum_{i=1}^{K_{g}}\mu_{1,i}+h_{Q_{1}}\sum_{i=1}^{K_{g}}\sqrt{1-\mu^{2}_{1,i}}\Bigg{)}$
=$=\Bigg{\{}\Bigg{(}h_{I_{1}}\sum_{i=1}^{K_{g}}\mu_{1,i}+h_{I_{1}}\sum_{i=1}^{K_{g}}\sqrt{1-\mu^{2}_{1,i}}\Bigg{)}^{2}$
$+\Bigg{(}h_{Q_{1}}\sum_{i=1}^{K_{g}}\mu_{1,i}+h_{Q_{1}}\sum_{i=1}^{K_{g}}\sqrt{1-\mu^{2}_{1,i}}\Bigg{)}^{2}\Bigg{\}}^{\frac{1}{2}}$
×$\times\arctan\!\!\left(\dfrac{h_{Q_{1}}\displaystyle\sum_{i=1}^{K_{g}}\mu_{1,i}+h_{Q_{1}}\sum_{i=1}^{K_{g}}\sqrt{1-\mu^{2}_{1,i}}}{h_{I_{1}}\displaystyle\sum_{i=1}^{K_{g}}\mu_{1,i}+h_{I_{1}}\sum_{i=1}^{K_{g}}\sqrt{1-\mu^{2}_{1,i}}}\right)\!\!.$ (12)
@@
$\phi^{opt}_{r,b}\!=\!\!\!\!\sum_{p\in\{h,g\}}\!\!\!\!\arctan\!\!\left(\!\!\frac{p_{Q_{1}}\displaystyle\sum_{i=1}^{K_{g}}\mu_{1,i}+p_{Q_{1}}\sum_{i=1}^{K_{g}}\sqrt{1-\mu^{2}_{1,i}}}{p_{I_{1}}\displaystyle\sum_{i=1}^{K_{g}}\mu_{1,i}+p_{I_{1}}\displaystyle\sum_{i=1}^{K_{g}}\sqrt{1-\mu^{2}_{1,i}}}\!\!\right)$ (13)
@@
$\hat{\gamma}_{j,k}^{r,b}=\zeta\Big{|}h_{U_{j},R_{r,b}}\Big{|}^{2}\Big{|}g_{R_{r,b},U_{k}}\Big{|}^{2}\Big{|}e^{-\delta}\Big{|}^{2}$ (15)

### v2

# Priority-aware grouping-based multihop routing scheme for RIS-assisted wireless networks

@@ Abstract
Reconfigurable intelligent surfaces (RISs) is a novel communication technology that has been recently presented as a potential candidate for beyond fifth-generation wireless communication networks. In this paper, we propose a priority-aware user traffic-dependent grouping-based multihop routing scheme for a RIS-assisted millimeter wave (mmWave) device-to-device (D2D) communication network with spatially correlated channels. Specifically, the proposed scheme exploits the priority of the users (based on their respective delay-constrained applications) and the aspect of spatial correlation in the narrowly spaced reflecting elements of the RISs. Here, based on the other users in the neighborhood, their respective traffic characteristics, and the already deployed RISs in the surroundings, we establish a multihop connection for information transfer from one of the users to its intended receiver. In this context, we take into account the impact of considering practical discrete phase shifts at the RIS patches instead of its ideal continuous counterpart. Moreover, we also claim and demonstrate that the existing classic least remaining distance (LRD)-based approach is not always the optimal solution. Finally, numerical results demonstrate the advantages of the proposed strategy and that it significantly outperforms the existing benchmark schemes in terms of system performance metrics such as data throughput, energy consumption, as well as energy efficiency.
@@ I Introduction
In recent years, due to applications such as the Internet of Things (IoT) and enhanced mobile broadband communication, wireless traffic has increased exponentially; it is projected to expand by more than five times in the $2023-2028$ period [1]. As a result, various technologies such as adaptive modulation, beamforming, and multiple access frameworks have been developed in the last few decades to support this overwhelming data traffic. The unifying theme behind all these advancements is to intelligently exploit and adapt to the randomly fluctuating wireless channel instead of having a control over the same. On the contrary, a new technology termed as the reconfigurable intelligent surfaces (RISs) is aimed at controlling the random nature of the channel for the purpose of its maximum utilization [2]. RISs essentially consist of arrays of reconfigurable passive elements embedded on a flat metasurface, which effectively ‘controls’ the channel instead of adapting to its varying nature [3]. These passive elements can be switched ON and OFF by adjusting the bias voltage. Moreover, without the need of any radio-frequency chains, they can reflect the incident signal in a desired direction. This results in reducing the implementation cost, which also enhances the system energy efficiency [4]. Furthermore, multiple variations of the RISs [5, 6, 7] have also opened new directions of research. By incorporating imperfect channel state information and hardware impairments, the authors in [5] investigates simultaneosuly transmitting and reflecting RIS-assisted non-orthogonal multiple access schemes. In [6], beamforming algorithms are proposed for enhancing the performance of reconfigurable holographic surfaces assisted cell-free networks. The work in [7] investigates active RIS-aided systems and analyzes its performance.
@@
In this context, RIS-assisted device-to-device (D2D) communication [8, 9, 10] forms a very relevant and interesting direction of research. The authors in [8] focus on uplink D2D-enabled cellular networks assisted by an RIS, where multiple D2D links share the same spectrum. In [9], the outage and rate performance of an RIS-assisted D2D communication underlaying cellular network is investigated by considering a generalized Nakagami-$m$ fading scenario. By using the concept of Markov service processes, the work in [10] analyzes the throughput of a delay-constrained RIS-assisted D2D communication scenario. Apart from system energy efficiency improvement by using RISs, the aspect of obtaining a high data rate can also be addressed by the usage of high frequency signals, for example, the millimeter waves (mmWaves) [11]. Note that, although mmWaves solve the problem of enhanced data rate, it has its own set of shortcomings such as significant propagation and high penetration losses. As a result, RIS assisted mmWave-based D2D network appears to be the solution for such scenarios, especially where the direct line of sight (LoS) link is of inadequate strength. To fully exploit the advantages of such networks, the aspect of strategic RIS placement is extremely crucial as they need to have clear LoS links with both the devices that intend to communicate with each other. Optimal RIS placement [12, 13, 14] can significantly extend the coverage of such networks with a minimum number of RISs, which effectively reduces both the hardware and installation expenditure.
@@
We further note that the RIS is essentially a passive device that simply reflects the incoming signal as well as the noise to its desired direction by tuning its parameter. Moreover, any given pair of users always communicate with each other for only a finite amount of time. Thus, augmenting the environment with a large number of RISs may result in unnecessary wastage of resources. A possible solution to this problem is to adopt a cooperative multihop approach [15, 16], i.e., apart from the RISs, the other intermediate users (IUs) present in the surroundings, depending on whether they are busy or idle, can be used as relays, namely, decode and forward (DF) or amplify and forward (AF). While AF relays have low processing cost, DF relays have high reliability as they forward only the decoded information portion to the next hop and not the entire received information-plus-noise mixture [17]. Thus, the option of having the idle IUs to act as DF relays seems beneficial. However, the role of RISs cannot be totally overlooked as in case of delay-constrained scenarios and/or idle IU unavailability, they can provide the aspect of guaranteed reliability [18]. Note that, the aspect of delay-constrained communication is very crucial, especially in 5G applications like ultra reliable low latency communication (URLLC) [19]. In this context, [20] proposes an energy-efficient adaptive topology-management system for sustaining network connectivity in resource-constrained wireless sensor networks (WSNs). Also, [21] investigates a fuzzy logic-based framework for quality-of-service aware routing protocol in wireless body area networks.
Accordingly, there exists some works, which investigate various aspects of routing in RIS-aided multihop networks. The work in [22] proposes a practical hybrid beamforming architecture for multi-hop multiuser RIS-empowered wireless THz communication networks. Energy-efficiency (EE) being one of the critical issues nowadays, [23] investigates the aspect of EE maximization of the RIS-assisted multihop networks.
@@
Moreover, as far as the structure of a RIS is concerned, certain aspects are critical to be considered and analyzed in this multihop framework. Firstly, the arrays of the reconfigurable elements are very narrowly spaced, which implies that the channels between the consecutive reflecting elements and the source/destination are spatially correlated [24, 25]. Hence, to avoid the huge channel estimation overhead in RIS-based systems, the works in [26, 27] propose grouping strategies where the nearby RIS elements are grouped together into smaller non-overlapping surfaces with identical phase shifts. Secondly, most existing works assume a continuous nature of the phase shifts that are offered by a RIS whereas in practical systems, they are actually discrete in nature [28]. Hence, these limited phase shifts do have a significant impact on the system performance. These aspects of a RIS cannot be overlooked if we are interested in investigating the practically feasible of a joint user and RIS based multihop framework. Furthermore, least remaining distance (LRD)-based approach is a standard procedure in literature to establish a multihop connection between two users, whether the sole objective is to reduce the effective distance between the two in each hop [29]. But, unlike in this case, this approach stands valid if and only if we do not take the channel conditions and the availability of the IUs into consideration.
@@ II-A Network Topology
Each RIS has $\mathcal{K}$ reflecting elements, which are effectively controlled to adjust both the amplitude and phase of the incident waveform [^1] Here we do not consider secondary reflections from the RISs. However, the proposed framework is general and it can be readily extended to the case where the aspect of secondary reflection is incorporated [31].. However, for the sake of simplicity and mathematical tractability, the amplitude factor is set to unity and it is only the phase that is tuned or optimized [26]. To reduce the channel estimation overhead, we employ a grouping strategy, where nearby $K_{g}$ reflecting elements are grouped together, such that all the elements belonging to the same group introduce identical phase shift [26]. Specifically, the $r$-th RIS $R_{r}$, consisting of $\mathcal{K}$ reflecting elements, is subdivided into $B$ non-overlapping surfaces $R_{r,b}$ $\forall$ $1\leq b\leq B$ such that $BK_{g}=\mathcal{K}$, where the number of partitions and the resulting sub-surfaces are determined a-priori [26]. Without any loss of generality, we consider that each group has an equal number of elements and a particular group can serve one service request at a time [32]. In this scenario, $R_{r,b}$ has two possible states: ON and OFF. If an element is in ON state, the phase of an incident signal can be changed to a desired direction and it cannot reflect while in OFF state. Finally, an example of the considered scenario is demonstrated in Fig. 1, where $S_{1}$ has the highest priority and the elements of the RIS are grouped into four non-overlapping surfaces, and the signal from $S_{1}$ reaches $D_{1}$ by using the path $S_{1}-U_{1}-R_{1,1}-U_{3}-U_{4}-D_{1}$.
@@ II-B Channel Model
Here $K_{p,q}$ is the Rician factor corresponding to the channel gain $h_{p,q}$ and $I_{0}(\cdot)$ denotes the zero-order modified Bessel function of the first kind. Moreover, note that direct $S-D$ path does not exist, i.e., $h_{S_{i},D_{i}}=0$ $\forall$ $i$.
@@ III Proposed strategy
In this section, we propose a priority-aware scheduling strategy for grouping-based RIS-assisted wireless networks. Here, we consider a delay-constrained scenario, where the data from $S_{i}$ must arrive at $D_{i}$ within the delay bound $T_{d_{i}}$. Firstly, we characterize the achievable data rate for a grouping-based RIS-aided scenario. Secondly, based on the channel condition and the number of coding bits being used at the RIS, we comment on the suitable group selection strategy. Thirdly, depending on the channel condition as well as the delay sensitivity of the application, a novel scheduling strategy is proposed for information transfer between any two IUs. Note that, we also comment on the appropriate choice of modulation scheme and appropriate next hop selection criteria. Finally, we provide an illustrative example of the proposed framework.
@@ III-A Achievable Rate Characterization of Grouping-based RIS
$h_{S_{i}/U_{j},R_{r,b}}$ $=\left(h_{I1}\sum\limits_{p=1}^{K_{g}}\mu_{1,p}+\sum\limits_{p=1}^{K_{g}}h_{Ip}\sqrt{1-\mu_{1,p}^{2}}\right)$
$+j\left(h_{Q1}\sum\limits_{p=1}^{K_{g}}\mu_{1,p}+\sum\limits_{p=1}^{K_{g}}h_{Qp}\sqrt{1-\mu_{1,p}^{2}}\right)$ (7)
@@
where $\zeta=\frac{P\rho_{\rm L}^{2}\left(d_{U_{j},R_{r}}d_{R_{r},U_{k}}\right)^{-\alpha}}{\sigma_{0}^{2}}$, $\phi_{r,b}$ is the associated phase shift provided at $R_{r,b}$, $T_{c}$ is the channel coherence time normalized to the number of slots, and $\sigma_{0}$ is the power of the zero mean additive white Gaussian noise (AWGN) at $U_{k}$. Moreover, the quantity $\frac{2}{T_{c}}$ denotes the feedback and processing delay. Note that unlike the conventional RIS-based approach, $\gamma_{j,k}^{r,b}$ does not involve a diagonal phase shift matrix of $K_{g}$ non-zero elements. On the contrary, all the elements of $R_{r,b}$ provide a common phase shift $\phi_{r,b}$ to the incoming signal. Hence, by rewriting the complex composite channels as $h_{U_{j},R_{r,b}}\!=\!\Big{|}h_{U_{j},R_{r,b}}\Big{|}e^{-j\theta_{h}}$ and $g_{R_{r,b},U_{k}}\!=\!\Big{|}g_{R_{r,b},U_{k}}\Big{|}e^{-j\theta_{g}}$, we obtain the optimal phase shift corresponding to $R_{r,b}$ as $\phi_{r,b}^{\rm opt}=\theta_{h}+\theta_{g}$ such that the maximum attainable SNR is given by (III-A). Finally, as we focus on the performance of idle IUs acting as DF relays, we consider an ideal scenario where we ignore the interference from other IUs; interference mitigation can be achieved through sophisticated signal processing and equalization techniques [36].
@@
=$\overset{(a)}{=}\frac{P\rho_{\rm L}^{2}\left(d_{U_{j},R_{r}}d_{R_{r},U_{k}}\right)^{-\alpha}}{\sigma_{0}^{2}}\left(\left(h_{I1}\sum\limits_{p=1}^{K_{g}}\mu_{1,p}+\sum\limits_{p=1}^{K_{g}}h_{Ip}\sqrt{1-\mu_{1,p}^{2}}\right)^{2}+\left(h_{Q1}\sum\limits_{p=1}^{K_{g}}\mu_{1,p}+\sum\limits_{p=1}^{K_{g}}h_{Qp}\sqrt{1-\mu_{1,p}^{2}}\right)^{2}\right)$
×$\times\left(\left(g_{I1}\sum\limits_{p=1}^{K_{g}}\mu_{1,p}+\sum\limits_{p=1}^{K_{g}}g_{Ii}\sqrt{1-\mu_{1,p}^{2}}\right)^{2}+\left(g_{Q1}\sum\limits_{p=1}^{K_{g}}\mu_{1,p}+\sum\limits_{p=1}^{K_{g}}g_{Qi}\sqrt{1-\mu_{1,p}^{2}}\right)^{2}\right),$ (10)
@@ III-B Group Selection Criteria at an RIS
$\!\!h_{U_{j},R_{r,b}}\!\!$ $=\Bigg{(}h_{I_{1}}\sum_{p=1}^{K_{g}}\mu_{1,p}+h_{I_{1}}\sum_{p=1}^{K_{g}}\sqrt{1-\mu^{2}_{1,p}}\Bigg{)}$
$+j\Bigg{(}h_{Q_{1}}\sum_{p=1}^{K_{g}}\mu_{1,p}+h_{Q_{1}}\sum_{p=1}^{K_{g}}\sqrt{1-\mu^{2}_{1,p}}\Bigg{)}$
=$=\Bigg{\{}\Bigg{(}h_{I_{1}}\sum_{p=1}^{K_{g}}\mu_{1,p}+h_{I_{1}}\sum_{p=1}^{K_{g}}\sqrt{1-\mu^{2}_{1,p}}\Bigg{)}^{2}$
$+\Bigg{(}h_{Q_{1}}\sum_{p=1}^{K_{g}}\mu_{1,p}+h_{Q_{1}}\sum_{p=1}^{K_{g}}\sqrt{1-\mu^{2}_{1,p}}\Bigg{)}^{2}\Bigg{\}}^{\frac{1}{2}}$
×$\times\arctan\!\!\left(\dfrac{h_{Q_{1}}\displaystyle\sum_{p=1}^{K_{g}}\mu_{1,p}+h_{Q_{1}}\sum_{p=1}^{K_{g}}\sqrt{1-\mu^{2}_{1,p}}}{h_{I_{1}}\displaystyle\sum_{p=1}^{K_{g}}\mu_{1,p}+h_{I_{1}}\sum_{p=1}^{K_{g}}\sqrt{1-\mu^{2}_{1,p}}}\right)\!\!.$ (12)
@@
$\phi^{opt}_{r,b}\!=\!\!\!\!\sum_{p\in\{h,g\}}\!\!\!\!\arctan\!\!\left(\!\!\frac{p_{Q_{1}}\displaystyle\sum_{p=1}^{K_{g}}\mu_{1,p}+p_{Q_{1}}\sum_{p=1}^{K_{g}}\sqrt{1-\mu^{2}_{1,p}}}{p_{I_{1}}\displaystyle\sum_{p=1}^{K_{g}}\mu_{1,p}+p_{I_{1}}\displaystyle\sum_{p=1}^{K_{g}}\sqrt{1-\mu^{2}_{1,p}}}\!\!\right)$ (13)
Based on the considered grouping framework, $\phi^{opt}_{r,b}$ is the optimal phase shift corresponding to all the elements of $R_{r,b}.$ Here, we state to differentiate our grouping framework from the one presented in [37]. This work considers a RIS-assisted aerial-terrestrial communication system, where the RIS is attached to a building, assisting the downlink communications of multiple UAV-user pairs. Here, the RIS is sub-divided into multiple groups, such that, one RIS group can serve one UAV-user pair. However, within a group, each reflecting element is allotted its own unique phase shift. On the contrary, in our proposed framework, all the elements belonging to the same group introduce identical phase shift, resulting in reduced channel estimation overhead.
@@
$\hat{\gamma}_{j,k}^{r,b}=\zeta\Big{|}h_{U_{j},R_{r,b}}\Big{|}^{2}\Big{|}g_{R_{r,b},U_{k}}\Big{|}^{2}\Big{|}e^{-j\delta}\Big{|}^{2}$ (15)